RC meeting and Run Plan, Tuesday Sept. 5, 2023

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time:at 3:00pm today

  • Location: Counting house conference room
  • ZoomGov Meeting Connection Information


Useful Links

Readiness for Beam

Hall readiness and lockup schedule (Zak Remele)

  • ZR: working on getting Jack's lock off the power supply; but beyond that it's standard preparation

Beamline readiness: beamline quads and harps, bcms, bpms, ion chambers (Ciprian Gal)

  • Daniel Moser: Will set up beam plans for reviving the target protection systems (the position locks and ramp rate). Will work with Bill Henry to coordinate and ensure the plan is ok. The discontinuities in the current readback are most likely due to the uneven sampling of the BCM readbacks (investigated last run and saw that discontinuities in the hall BCMs were not reflected in the injector line BCMs).
  • Ion chambers are still being checked out by SSG.

Subsystems Checklist

For each subsystem we ask that the responsible person go over the following points:

  1. The overall status of your subsystem
  2. Will your subsystem be ready well before Hall lock-up on Friday or at the latest by Monday morning?
  3. Does your system being ready depend on your team only or are outside resources required?


  • HCal (Jiwan Poudel)
    • few channels with problems (probably cabling) and expect to recover them ahead of closeout on Friday
  • Bigbite (Arun Tadepalli)
    • already took some cosmic data and things look ok
  • Target cell status(Arun/Gordon)
    • Plan in place to complete work ahead of Friday
  • Target laser status (alignment, polarization, power etc)
    • work ongoing
  • Counting House Alarms (Bill Henry)
    • Bill will test the alarms and instructions for shift crew response in the wiki
  • DAQ (Alex Camsonne)
    • SBS beamline crate recovery ongoing but expect to have it done by Friday
    • there seems to be a sitewide helicity signal issue; will talk to Paul King
  • GEMs (Holly Szumila-Vance)
    • Sean: SBS GEMs require debugging (but they should not hold off closing the hall)
  • Target magnets and measurements (Bogdan W. and Bill Henry)
    • Plan to do field measurements Wed morning
  • Moller polarimeter (Don Jones)
    • Target motion not working (Ellen looking into it)
    • we want to do at least one measurement at the beginning; all systems that could be checked now (besides the target) work as expected
  • SBS magnet and Bigbite magnets ramp and hold?
    • Power supply is locked by Jack and paperwork is being done to take it off and test them
  • Corrector magnets:
    • need to be tested by Ellen; CG provided max current
  • Hodoscopes:
    • need update

Plans and Procedures

Are required plans and procedures in place?

  • Is there a clear and up to date procedure for bring beam into the hall for the first time?(Eg. [1])
  • Are the previous target protections still in place? Do we need to remind MCC to re-establish them?
    • Reminder: software GUIs with target position protection using BPMs engaged and turns beam off if it wanders +/1mm from set position. Must go to carbon target or target out to tune if beam trips due to this.
    • Hall A current ramp protection has maximum ramp rate and does what??? if it goes over a threshold above the set ramp rate.
  • Are the existing written procedures for re-establishing beam after being down for a period of time still up to date. Are the thresholds triggering the re-establishment procedure still valid?
  • Can we continue using the same spreadsheet?
  • Is there a clear experimental plan for the RC to accomplish for the first week of running? (Andrew Puckett and Bogdan Wojtsekhowski)
    • When will we do polarimetry? Are there optics runs we need to take? What calibrations need to be done? Any sieve in data?